# In-context Learning in Presence of Spurious Correlations

### v1

@@ 3 In-context learning based on a diverse set of tasks
In Section 2, we showed that it is possible to obtain a good in-context learner for a given task, but it fails to generalize to tasks with different spurious features. A better in-context learner should detect spurious features from context and make predictions without employing them. In this section, we explore the possibility of obtaining such a learner by training on a diverse set of ICL tasks. Since there exist few suitable datasets, we synthesize binary classification tasks with a single binary spurious feature, aiming to capture “structure” present in existing datasets. In short, given a standard binary classification task, say cat vs dog classification, for a sampled minority of cats we overwrite some of their features with those of random dogs. Similarly, we do an analogous operation for a sampled minority of dogs. This way some cats share dog features and vice versa. To create a diverse pool of in-context learning instances, we vary the two classes and the subset of grafted features. Please refer to Figure 6 for an illustration of this grafting operation.
More concretely, we consider the iNaturalist dataset (Van Horn et al., 2018), which contains images from 5,089 natural fine-grained categories and filter out categories that have less than 500 images. For testing purposes, from remaining 239 categories we set apart categories that belong to the supercategories Amphibia and Mammalia, along with 10% of random categories. We denote the set of these 48 categories as $\mathcal{C}_{\mathrm{ood}}$, and the set of remaining 191 categories as $\mathcal{C}_{\mathrm{id}}$, which we use to create in-context learning instances for training. For each category in $\mathcal{C}_{\mathrm{id}}$, we hold out half of the examples as in-distribution validation set. To generate a single in-context learning instance, we sample two distinct classes from $\mathcal{C}_{\mathrm{id}}$ randomly and sample $n/2$ images from the training split of each class uniformly at random without replacement. We then do the grafting operation, setting minority group ratio within each class to 10%. We select the grafted features randomly, by first picking subset size $k$ uniformly at random from 0 to 199, and then sampling a random subset of embedding dimensions of size $k$. With this we get $n$ examples that form the context part of the instance. Abandoning the naive approach and focusing on the proposed one, for each class we sample $n/2$ queries from the remaining examples uniformly at random with replacement and do the grafting operation with 50% minority group ratio.
Following the experiments in Section 2, we train the same transformer with the proposed approach on 4M ICL instances with $n=400$ context examples. We use the same optimizer and sweep the learning rate in the same range, selecting the best value based on the average minority-group accuracy (defined exactly in Appendix B) on instances where both categories belong to $\mathcal{C}_{\mathrm{ood}}$ and thus were not observed during training. The results presented in Figure 7 indicate a major difference compared to the results in the single-task regime – namely, the proposed approach learns to do in-context learning to some extent without permuting embedding dimensions. As expected, we see much better performance with permuted embedding dimensions. Notably, comparing majority-group and minority-group accuracies of the proposed approach with permutations, we see almost no sign of reliance on spurious features.
Promoting emergence of induction heads.
In-context learning ability has been linked to induction heads, which are specific type of circuits found within large language models that implement the operation of looking back over the sequence for finding previous instances of the current token and copying what comes after that (Olsson et al., 2022). Inspired by this, we propose a data preparation technique that promotes learning of induction heads. With probability $p$, we replace each intermediate query independently with a random example from the proceeding part of the context. Note that this type of “hinting” is not possible in the naive approach and is enabled by the introduction of intermediate queries. In all experiments with this technique enabled, we just set $p=0.25$. We observed that training of typical runs escapes the initial loss plateau faster with this technique (in about 3k iterations compared instead of about 10k iterations). Moreover, we see modest performance gains in iNaturalist experiments (see Figure 7, where +I stands for this technique).
Figure 8: Minority-group accuracy on the OOD test set of iNaturalist for the best proposed approach with or without passing group information as input. Shaded regions show standard deviation across 5 training runs.
Figure 9: Minority-group accuracy on the OOD test set of iNaturalist for the best variant of proposed approach and conventional methods such as 1-NN, ERM, and GroupDRO. Shaded regions show standard deviation across 5 training runs.
Passing example groups as input.
In contrast to the findings in the single-task setting of Section 2, we observed that setting $\tilde{y}_{i}$ to represent group improves the proposed approach, even on top of permitting input dimensions and promoting induction heads. One case of this is presented in Figure 9, while more cases can be found in the complete results presented in Appendix B. For brevity, we mark passing groups as inputs with +G in figures and tables.
Comparison with conventional learning algorithms.
Similar to the experiments in Section 2, we compare the best variant of the proposed approach (G + P + I) to 1-NN, ERM, and GroupDRO. Results presented in Figure 9 show that the learned algorithm is on-par with or outperforms the baselines starting at context length 32. The results at context lengths below 20 are not as informative, because the way we implemented the grafting operation implies that no examples are grafted when there are less than 10 examples in a class.
Figure 10: Worst-group test accuracy on Waterbirds for the best variant of proposed approach trained on iNaturalist and for methods such as 1-NN, ERM, and GroupDRO. Shaded regions show standard deviation across 5 training runs.
Figure 11: Worst-group test accuracy on Waterbirds-severe for the best variant of proposed approach trained on iNaturalist and for methods such as 1-NN, ERM, and GroupDRO. Shaded regions show standard deviation across 5 training runs.
Generality of the learned algorithm.
To test the generality of the learned algorithm, we report evaluation results on Waterbirds (Figure 11) and Waterbirds-severe (Figure 11). We see that the learned algorithm outperforms baselines on Waterbirds and is as good as we got by training on Waterbirds itself. However, the learned algorithm fails completely on Waterbirds-severe, while the baselines give meaningful results starting at context length 32. We hypothesize that the challenge posed by the spurious features in Waterbirds-severe is significantly more severe compared to that in iNaturalist. By varying the norm of the added background vector, we interpolate between Waterbirds and Waterbirds-severe, and we see good generalization until the norm of the added vector is $\approx$40of the average embedding norm (see Figure 12).
4 Related work
In this section, we discuss more related work in addition to the ones discussed earlier.
In-weights vs in-context learning.
We observe two modes of learner behavior in our experiments. In the first mode, the learner acts like a standard supervised classifier, ignoring context examples. This mode appears when training on ICL instance of a single task without permuting input embedding dimensions. In the second mode, the learner does proper in-context learning. Our experiments indicate that both permuting embedding dimensions and increasing the number of training tasks are reliable ways of steering the model towards the in-context learning mode. The former is akin to the method of randomly projecting inputs proposed by Kirsch et al. (2022) for obtaining general-purpose in-context classifiers. Prior work has made a distinction between these two models of learning, naming them in-weights and in-context learning. In particular, Chan et al. (2022) demonstrate that certain distributional properties of data, such as long-tail of class frequencies and bursty distribution of context example classes, can promote in-context learning when meta-training on few-shot classification instances. Singh et al. (2024) show that in-context learning behavior is not persistent and decays away with overtraining, indicating a trade-off between in-weights and in-context learning mechanisms. Moreover, they find that this in-context learning skill decay can be prevented by applying weight decay of embeddings and MLP layers, slowing down in-weight learning. Anand et al. (2024) make similar observations about these two modes of learning and propose active forgetting of token embeddings as an effective way of steering towards the in-context learning mode.
Many shot ICL.
One ancillary finding of this work is that transformers can be trained to do in-context learning of visual classification tasks when good image embeddings are provided. This is remarkable because the input dimensionality we considered is much higher than what was considered in the pioneering works of Garg et al. (2022) and Akyürek et al. (2022) (784 vs 20). Furthermore, we observe predictable performance gains from longer context sizes. The number of “shots” we consider (up to 512 examples) is well beyond what is typically considered in ICL works (up to a few dozen of examples). Our findings are complementary to those of Agarwal et al. (2024)Jiang et al. (2024), and Li et al. (2024) who find that multimodal large language models, such as Gemini-1.5 Pro and GPT-4o, can benefit from large number of in-context demonstrations (up to 1000 demonstrations).
In-context learning for out-of-distribution generalization.
Closest to our work are the works that propose to make use of in-context learning for out-of-distribution generalization. Han et al. (2023) test multimodal large language models (MLLMs) on a variety of visual classification tasks. They propose to leverage in-context learning abilities of MLLMs to improve performance on specialized domains and on tasks with significant corruptions. However, they only consider the case where both context examples and query are from the target domain. Zhang et al. (2024) make similar observations, but additionally study robustness of in-context learning to distribution shifts, such as domain shifts, label shifts, and spurious correlations. They find that in-context learning is highly susceptible to label shifts and presence of spurious correlations. Finally, Gupta et al. (2024) propose to address the problem of domain generalization (Muandet et al., 2013) by training an in-context learner that can take examples from a domain/environment and adapt to that domain in-context.
5 Discussion and conclusion
Spurious correlations can appear in various domains, often in complex and unexpected ways. For example, a medical device might make a false positive or worse, a false negative prediction due to the specifics of the patient and training data. Given the increasing reliance on in-context classification, particularly with the advancements in large language models, understanding and mitigating the negative effects of spurious correlations is important for ensuring the reliability and safety of these models.
We showed that it is possible to train an effective in-context learner tailored to a particular classification task with spurious features. We did this by introduce two key techniques: (a) permuting input embedding dimensions and (b) forming ICL sequences with intermediate queries simulating distribution shift. We provided evidence that the learned algorithm is highly competitive on the task it was trained on. However, we found that while it generalizes to other tasks without spurious features, it does not work for tasks with other spurious features. Understanding this failure mechanistically and exploring techniques for enabling better generalization are key future research directions.
We next explored training on synthetic ICL instances of diverse tasks and showed that it is possible to obtain an in-context learner that generalizes to unseen tasks, even with different data generating processes. We established the usefulness of two more techniques: (c) passing example groups as input and (d) promoting learning of induction heads by occasionally querying past context examples. We believe there is a room for improving in-context learning via improved strategies of choosing intermediate queries and possibly optimizing worst-group loss. Understanding why the learned algorithm fails under extreme distribution shifts and why variants with permutations fail more (see Figure 12) is an interesting question to explore. Another interesting direction to explore is to find out what exact algorithm is learned in the process of training on diverse tasks. Based on the results presented in this work, we conclude that the learned algorithm is neither 1-NN, ERM, or GroupDRO.
Our work has several limitations. First of all, training a transformer-based in-context learner with high-dimensional image embeddings is computationally costly (see Appendix A for information on compute resources), although it is faster than the baselines at inference. For this reason, we did not explore more datasets and pretrained image embeddings. We believe main conclusions of our work will be unchanged. Second, we experimented with only one model size, width, and depth. Larger models might behave differently (Wei et al., 2023). Third, in our iNaturalist experiments, we considered only one “type” of spurious features. It is likely that this choice has significant effect on the learned algorithm and its generality. Future research should explore more ways of synthesizing spurious features and consider varying severity of the challenge posed by spurious features. The latter can be done by considering multiple spurious features, introducing label imbalance, varying magnitude of spurious correlations, and varying the margin spurious features provide. Finally, we acknowledge that the proposed approach requires spurious feature annotations, which is typically costly to obtain. However, they are required only at training time, not at inference, As we have shown, this limitation can be addressed by creating synthetic data, in which case spurious annotations are readily available.
Acknowledgments
We would like to thank anonymous reviewers for valuable feedback. Samvel Karapetyan’s work is supported by Yandex Armenia Scholarship.
Appendix A Further experimental details
Baselines.
For empirical risk minimization as a baseline, we tune 2 hyperparameters: learning rate (0.01 or 0.001) and number of epochs (100 or 200). For GroupDRO we additionally tune its parameter that controls adaptiveness of group weights (0.01, 0.1, or 1) and we also try an optional strong L2 regularization (1.0 weight decay), as it has been observed to be useful for small datasets [Sagawa* et al., 2020].
Transformer-based methods.
In all transformer-based approaches, we train a causal decoder-only GPT-J transformer with 80M parameters that has 6 transformer layers with 8 multi-head attention, 768 model dimensionality, and 3072 hidden dimensionality. When training on iNaturalist, we add a layer normalization [Ba et al., 2016] on transformer input, as we expect input norms to change when we evaluate on Waterbirds-based datasets. The transformer input sequence in the proposed approach consists of 3 types of tokens: context image embeddings, query image embeddings, and label/group annotations. While the network can rely on positions and content to distinguish image embeddings from annotations, we found it to be helpful to encode token types explicitly. We do this by setting the first 3 dimensions of a token to be a one-hot vector representing token type (context image embedding, query image embedding, or annotation). When permuting dimensions, we do the permutation before encoding token types to keep the location of token-type information consistent. In our preliminary experiments and development, we used $n=128$ context length. Apart from improved performance, we did not observe significant qualitative differences when we switched to larger context lengths for final experiments.
Evaluation and model selection.
For all transformer-based approaches and baselines, we do a grid search to find the best combination of hyperparameters. In particular, we train each configuration with 5 different random seeds and selected one with the highest average test performance. Importantly, for baseline methods model selection is done for each context length independently, while for transformer-based methods model selection is done once with respect to the test performance at maximum context length observed during training. All evaluations are done on 8192 sequences, where the first $n$ examples are sampled from the corresponding train set while the query is sampled from the test set with a balanced group distribution. Finally, even when training transformers on permuted image embeddings, we do not apply permutations during evaluation. In all figures throughout this work, shaded regions show standard deviation across the 5 training runs.
Note that the most principled model selection approach would be selecting models based on a metric calculated on a dataset similar to the training set (e.g., a held-out part of training set), rather than the test set. For example, in the case of experiments on Waterbirds or Waterbirds-severe, the principled approach would be to select based on performance on sequences where the context part is sampled from the training set, while the final query is sampled from a held-out validation set with balanced group distribution. We tried this way of model selection and did not observe significant changes. In the case of experiments on iNaturalist, the principled approach would be to select based on performance on sequences where the context part is sampled from the training set, while the final query is sampled from the hold-out part the training set. We observed that this in-distribution metric is always around 99.5%-100%, and can be non-informative for model selection. This is a typical scenario in OOD generalization (see for example  [Gulrajani and Lopez-Paz, 2021] or [Wenzel et al., 2022]).
Definitions of metrics.
Given a set of predictions on Waterbirds or Waterbirds-severe, worst-group accuracy is defined as the lowest accuracy of predictions among the 4 groups. Note that worst-group accuracy is not applicable to iNaturalist, as different ICL sequences correspond to different classification tasks and hence form different groups. For this reason, we introduce minority-group and majority-group accuracies. Given a triplet $(C,q,\widehat{y})$, where $C$ is a context, $q$ is query, and $\widehat{y}$ is a prediction on $q$, we call $\widehat{y}$ a minority (majority) prediction, if $q$ is among the least (most) represented group(s) of the context $C$. Given a list of triplets $(C,q,\widehat{y})$, we define minority (majority) group accuracy as the accuracy among minority (majority) predictions.
Compute resources.
Appendix B Additional results
@@
In addition to the figures presented in the main text, here we provide the exact experimental resources for multiple transformer-based and baseline approaches, some of which were not included in the main text due to space constraints. Recall that +P means permuting input dimensions, +I means promoting learning of induction heads, and +G means passing example groups as input to in-context learning transformers.
@@
Figure 13: Majority-group and worst-group test accuracies on Waterbirds-severe as a function of context size for the naive approach with a single modification of making the last example (query) group-balanced. Shaded regions show standard deviation across 5 training runs. As expected, at intermediate context lengths this method performs similar to the naive approach, but is much better at the training context length.
@@
Table 1 presents worst-group accuracies on the test set of Waterbirds for 3 sets of approaches: (a) in-context learners trained on Waterbirds itself, (b) in-context learners trained on iNaturalist, and (c) baselines. Similarly, Table 2 presents worst-group accuracies on the test set of Waterbirds-severe for 3 sets of approaches: (a) in-context learners trained on Waterbirds-severe itself, (b) in-context learners trained on iNaturalist, and (c) baselines. As RoPE-based transformers are not good at length extrapolation [Press et al., 2021], we do not attempt evaluating models trained on iNaturalist with context size 400 on 512-long sequences of Waterbirds or Waterbirds-severe. Finally, Table 3 presents minority-group accuracy on out-of-distribution classes of iNaturalist for two sets of approaches: (a) in-context learners trained on iNaturalist itself and (b) baselines.
@@
To further verify our main findings presented in Section 2, we conduct experiments on another popular visual classification tasks CelebA [Liu et al., 2015]. In CelebA, the task is to classify blond vs non-blond persons, with sex being a spuriously correlated variable. Notably, the spurious correlation is asymmetric, in the sense that blond and non-blond women are almost equally represented, while blond men are much less represented compared to non-blond men. We follow the design of Waterbirds experiments in our CelebA experiments, with the only difference that we set the group distribution of context examples to $(0.25,0.25,0.05,0.45)$, where group 0 are non-blond men, group 1 are non-blond women, group 2 are blond men, and group 4 are blond women. Table 4 presents worst-group accuracies on the test set of CelebA for 2 sets of approaches: (a) in-context learners trained on CelebA itself and (b) baselines algorithms. As in our Waterbirds experiments, we see that it is essential to permute input embeddings and to form ICL sequences in the proposed fashion. Unlike Waterbirds, comparing “Proposed + P” with “Proposed + P + G” we see that providing spurious annotations in-context provides significant gains. Figure 4 demonstrates that both of these approaches outperform 1-NN, ERM, and GroupDRO.
@@
Table 1: Complete results on Waterbirds. Reported numbers are average worst-group test accuracies, along with the their standard deviation. The top half of in-context learners were trained on Waterbirds itself, while the ones in the bottom half were training on iNaturalist.
@@
Table 2: Complete results on Waterbirds-severe. Reported numbers are average worst-group test accuracies, along with the their standard deviation. The top half of in-context learners were trained on Waterbirds-severe itself, while the ones in the bottom half were training on iNaturalist.
@@
Table 3: Complete results on iNaturalist. Reported numbers are average minority-group accuracies on the OOD test set of iNaturalist, along with the their standard deviation.
@@
Table 4: Complete results on CelebA. Reported numbers are average worst-group test accuracies, along with the their standard deviation. All in-context learning were train on CelebA itself.

### v2

@@ Appendix B Additional illustrations
Appendix C Additional results
@@
In addition to the figures presented in the main text, here we provide the exact experimental results for multiple transformer-based and baseline approaches, some of which were not included in the main text due to space constraints. Recall that +P means permuting input dimensions, +I means promoting learning of induction heads, and +G means passing example groups as input to in-context learning transformers.
Table˜1 presents worst-group accuracies on the test set of Waterbirds for 3 sets of approaches: (a) in-context learners trained on Waterbirds itself, (b) in-context learners trained on iNaturalist, and (c) baselines. Similarly, Table˜2 presents worst-group accuracies on the test set of Waterbirds-severe for 3 sets of approaches: (a) in-context learners trained on Waterbirds-severe itself, (b) in-context learners trained on iNaturalist, and (c) baselines. As RoPE-based transformers are not good at length extrapolation (Press et al., 2021), we do not attempt evaluating models trained on iNaturalist with context size 400 on 512-long sequences of Waterbirds or Waterbirds-severe. Finally, Table˜3 presents minority-group accuracy on out-of-distribution classes of iNaturalist for two sets of approaches: (a) in-context learners trained on iNaturalist itself and (b) baselines.
Figure˜17 presents majority-group accuracies on Waterbirds and Waterbirds-severe for the naive and proposed methods. Figure˜18 presents majority-group accuracies on Waterbirds and Waterbirds-severe for the proposed approach and conventional methods.
(a) Waterbirds
(b) Waterbirds-severe
@@
Figure 17: Majority-group accuracies on Waterbirds and Waterbirds-severe as a function of context size for the naive and proposed approaches with or without permuting input dimensions. Shaded regions show standard deviation across 5 training runs.
@@
Figure˜19 demonstrates that it is important to simulate distribution shift at all intermediate query locations, not just for the main query. This makes sure that the learned network can make robust predictions for all intermediate context sizes.
Figure 19: Majority-group and worst-group test accuracies on Waterbirds-severe as a function of context size for the naive approach with a single modification of making the last example (query) group-balanced. Shaded regions show standard deviation across 5 training runs. As expected, at intermediate context lengths this method performs similarly to the naive approach, but is much better at the training context length.
(a) CelebA
(b) MultiNLI
Figure 20: Worst-group test accuracies on CelebA and MultiNLI for the proposed approach and conventional methods such as 1-NN, ERM, and GroupDRO. Shaded regions show standard deviation across 5 training runs.
@@
To further verify our main findings presented in Section˜2, we conduct experiments on CelebA (Liu et al., 2015). Here the task is to classify blond vs non-blond persons from face images, with sex being a spuriously correlated variable. Notably, the spurious correlation is asymmetric, in the sense that blond and non-blond women are almost equally represented, while blond men are much less represented compared to non-blond men. We follow the design of Waterbirds experiments in our CelebA experiments, with the only difference that we set the group distribution of context examples to $(0.25,0.25,0.05,0.45)$, where group 0 are non-blond men, group 1 are non-blond women, group 2 are blond men, and group 4 are blond women. Table˜4 presents worst-group accuracies on the test set of CelebA for 2 sets of approaches: (a) in-context learners trained on CelebA itself and (b) baseline algorithms. As in our Waterbirds experiments, we see that it is essential to permute input embeddings and to form ICL sequences in the proposed fashion. Unlike Waterbirds, comparing “Proposed + P” with “Proposed + P + G” we see that providing spurious annotations in-context provides significant gains. Figure˜20(a) demonstrates that both of these approaches outperform 1-NN, ERM, and GroupDRO.
Experiments on MultiNLI.
Beyond visual classification tasks, we also conduct experiments on a natural language inference (NLI) task. In particular, we consider the MultiNLI dataset (Williams et al., 2018b), where given a pair of sentences, a premise and a hypothesis, the task is to determine whether the hypothesis is entailed by, neutral with, or contradicts the premise. Prior work has observed that there is a spurious correlation between contradictions and the presence of the negation words nobody, no, never, and nothing (Gururangan et al., 2018). For our experiments, we frame a binary classification task entails or neutral vs. contradicts, with a binary spurious feature, which is 1 if and only if the hypothesis has one of the four negation words. We follow the design of Waterbirds experiments and set the group the group distribution of context examples to $(0.45,0.05,0.05,0.45)$. Table˜6 presents worst-group accuracies on the test set of MultiNLI for 2 sets of approaches: (a) in-context learners trained on CelebA itself and (b) baseline algorithms. Similar to Waterbirds and CelebA results, we see that the proposed approach of intermediate queries with balance group distribution helps to decrease the reliance on the spurious feature. However, in contrast to Waterbirds and CelebA experiments, permuting input embeddings has smaller effect on preventing in-weight learning. Figure˜20(b) demonstrates the "Proposed + P" method outperforms 1-NN, ERM, and GroupDRO.
Experiments with a larger network.
To verify that our findings generalize to larger models, we repeat CelebA experiments but with a transformer architecture of 12 layers with 12 multi-head attention (instead of 6 layers with 8 multi-head attention). Due to memory increase, we decrease the batch size from 24 to 8. Besides these two changes, we keep all other experimental details the same. The complete results presented in Table˜5 are qualitatively the same compared to the smaller network case (Table˜4), with the difference that the results of transformer-based entries are lower. Furthermore, the standard deviation of the +P+P approaches is significantly higher, indicating difficulties in optimization. We hypothesize that this is due to reusing learning rate and training length that were tuned for the smaller network with 3 times larger batch size.
Figure 21: Majority-group and worst-group test accuracies on Waterbirds as a function of context size for a “Proposed + P” run evaluated on ICL sequences with randomly rotated input embeddings. Largely unchanged evaluation results fail to confirm that there is any data leakage when input embeddings are permuted during training.
On data leakage in single task regime.
In the single task setting of Section˜2, there is a potential for data leakage, not in the sense that individual examples might be leaked (we always evaluate on unseen examples), but in the sense that the learner effectively observes more data from the single task than its context length at evaluation. Indeed, when we do not permute input embeddings, we observe task memorization (i.e., data leakage) and the model does very well at evaluation with even close to empty context. To verify that there is no data leakage when we enable permuting input embeddings (+P), we take one of the “Proposed + P” runs trained on Waterbirds and evaluate it on ICL sequences where input embeddings of each sequence are rotated with a random rotation matrix. As the set of permutation matrices is a measure-zero subset of general rotation matrices, we expect that in case of data leakage we would observe degraded performance, as the model would be expecting randomly permuted embeddings of some memorized embedding space. In results presented in Figure˜21, we see that under this new evaluation the results are the same (up to statistical noise), failing to confirm that there is any data leakage when input embeddings are permuted during training. Finally, note that data leakage is not a concern in the multiple task setting of Section˜3, because we evaluate on either unseen categories of iNaturalist or on unseen tasks such as Waterbirds and Waterbirds-severe.
Figure 22: Majority-group and worst-group test accuracies on Waterbirds as a function of context size for the naive approach trained and evaluated on group-balanced contexts. The training is done on ICL sequences with 128 context examples. Shaded regions show standard deviation across 5 training runs.
Experiments with group-balanced contexts.
As noted in Section˜5, the proposed approach of training an in-context learner requires spurious annotations. Given access to spurious annotations, one can simply train an in-context learner on sequences with balanced groups. While in-context learners obtained this way will not be useful for new tasks for which we do not have spurious annotations (and thus cannot form group-balanced contexts), it is still useful to compare how well this approach does in the single task setting of Section˜2. For this end, we train in-context learners on balanced-group sequences consisting of 128 Waterbirds examples. This way each group is represented with 32 context examples. Note that in our main Waterbirds experiments with 512 context examples but group-imbalanced contexts, the minority groups are represented with even fewer, 25 examples. As the group-balanced sampling context breaks the correlation between the label and spurious feature, we only consider the naive approach of forming ICL sequences (Figure˜2(a)). The results presented in Figure˜22 show that, as expected, group-balanced sampling improves worst-group accuracy. The naive approach, which again leads to in-weights learning, reaches 86.08 $\pm$ 1.87 worst-group accuracy with 128 group-balanced context examples, compared to 84.82 $\pm$ 1.26 worst-group accuracy on 512 group-imbalanced context examples (see Table˜1). This positive effect of downsampling has been also observed in standard (not in-context) training settings (Nagarajan et al., 2021; Menon et al., 2021; Idrissi et al., 2022). Furthermore, we again see that the proposed technique of permuting embedding dimensions induces strong in-context learning and reaches 90.44 $\pm$ 1.10 worst-group accuracy with 128 group-balanced context examples. This is just a bit below the 91.95 $\pm$ 1.20 worst-group accuracy we get training “Proposed + P” on 512 group-imbalanced context examples (see Table˜1).
@@
Table 1: Complete results on Waterbirds. Reported numbers are average worst-group test accuracies, along with their standard deviation. The top half of in-context learners were trained on Waterbirds itself, while the ones in the bottom half were training on iNaturalist.
@@
Table 2: Complete results on Waterbirds-severe. Reported numbers are average worst-group test accuracies, along with their standard deviation. The top half of in-context learners were trained on Waterbirds-severe itself, while the ones in the bottom half were training on iNaturalist.
@@
Table 3: Complete results on iNaturalist. Reported numbers are average minority-group accuracies on the OOD test set of iNaturalist, along with their standard deviation.
@@
Table 4: Complete results on CelebA. Reported numbers are average worst-group test accuracies, along with their standard deviation. All in-context learners were trained on CelebA itself.
@@
Table 5: Complete results on CelebA, but with larger network of 120m parameters, consisting of 12 layers (instead of 6 layers) with 12 multi-head attention (instead of 8 heads). Reported numbers are average worst-group test accuracies, along with their standard deviation. All in-context learners were trained on CelebA itself.
Method / Context size 4 8 16 32 64 128 256 512
Naive
24.43
21.79
23.68
23.02
23.99
23.18
22.62
20.22
Naive + P
21.34
14.64
12.56
13.35
13.76
15.73
17.69
21.67
Proposed
22.90
20.86
23.11
21.93
23.35
21.82
21.75
19.54
Proposed + P
35.13
31.54
30.89
35.19
41.63
47.81
51.60
55.53
Proposed + G
23.08
20.89
22.74
21.73
22.70
21.24
21.50
18.90
Proposed + G + P
31.44
32.09
36.54
43.67
49.62
54.77
58.75
61.74
1-NN
35.87
37.63
36.08
37.86
38.28
36.40
36.90
37.81
ERM
30.70
28.93
26.37
30.75
34.29
38.64
45.18
49.92
GroupDRO
35.32
34.64
30.24
37.49
47.11
54.56
56.11
61.47
Table 6: Complete results on MultiNLI. Reported numbers are average worst-group test accuracies, along with their standard deviation. All in-context learners were trained on MultiNLI itself.
Method / Context size 2 4 8 16 32 64 128 256 512
Naive
84.15
84.58
84.74
84.13
83.92
85.20
84.38
84.30
84.05
Naive + P
85.30
87.29
87.43
87.56
87.50
88.05
87.35
87.99
87.54
Proposed
86.62
86.02
87.45
87.54
86.99
86.51
86.11
87.35
86.89
Proposed + P
85.84
87.22
87.43
87.75
87.70
87.86
87.31
87.56
87.73
Proposed + G
87.15
86.23
87.23
87.37
87.02
86.51
86.47
87.69
87.50
Proposed + G + P
85.41
87.32
87.57
87.81
87.74
87.97
87.56
87.78
87.68
1-NN
64.49
82.41
85.84
86.39
86.01
86.04
86.44
85.90
86.53
ERM
76.40
81.71
85.10
86.82
86.76
87.30
87.04
87.10
87.55
GroupDRO
76.54
81.56
85.45
86.75
86.55
87.16
87.14
87.10
87.23